# Generative design of functional organic molecules for terahertz radiation detection

### v1

@@ Abstract
Plasmonic nanocavities are molecule-nanoparticle junctions that offer a promising approach to upconvert terahertz radiation into visible or near-infrared light, enabling nanoscale detection at room temperature. However, the identification of molecules with strong terahertz-to-visible upconversion efficiency is limited by the availability of suitable compounds in commercial databases. Here, we employ the generative autoregressive deep neural network, G-SchNet, to perform property-driven design of novel monothiolated molecules tailored for terahertz radiation detection. To design functional organic molecules, we iteratively bias G-SchNet to drive molecular generation towards highly active and synthesizable molecules based on machine learning-based property predictors, including molecular fingerprints and state-of-the-art neural networks. We study the reliability of these property predictors for generated molecules and analyze the chemical space and properties of generated molecules to identify trends in activity. Finally, we filter generated molecules and plan retrosynthetic routes from commercially available reactants to identify promising novel compounds and their most active vibrational modes in terahertz-to-visible upconversion.
@@ I Introduction
Terahertz (THz) radiation has applications in numerous fields, including medical diagnostics, security screening, communications, and astronomy [1, 2]. The 1–30 frequency range is often referred to as the terahertz gap, as the development of both powerful and affordable sources, and efficient wideband detectors, has been challenging for traditional electronics.
@@
The enhancement of electronic fields in plasmonic nanocavities can be utilized in molecular optomechanical devices which convert terahertz radiation to visible or near-infrared light [3, 4], enabling nanoscale, room temperature detection of terahertz radiation. To enhance the light-matter interaction, molecules are typically placed between two metallic nanoantennas [4, 5, 6]. One of the two antennas focuses terahertz radiation at the design frequency over the molecular sample volume to enhance the absorption of terahertz radiation via the surface-enhanced infrared absorption [7] mechanism. The second optical antenna confines visible or near-infrared light to volumes below 100 , which induces surface-enhanced Raman scattering [8] of molecules within the plasmonic nanocavity. Absorption of terahertz radiation by molecules within the nanocavity results in the vibrational excitation of a specific normal mode, which leads to an increase in the measured Raman anti-Stokes intensity of the same normal mode, similar to resonant sum-frequency generation spectroscopy [9]. For centrosymmetric molecules, simultaneous activity in absorption and Raman scattering is not possible. Even in asymmetric molecules, it is rare to have vibrational modes that can efficiently upconvert the terahertz radiation signal, which makes it necessary to use computational tools to quickly identify good candidate molecules and their active vibrational modes [10, 11].
@@
Machine learning (ML) methods can facilitate the design and discovery of new functional materials by enabling the fast computational screening of large structural databases [12, 13, 14]. ML-based screening has previously been used to identify promising candidates for terahertz radiation detection from commercially available compound databases [10]. However, a drawback of this approach was that there was a limited search pool of molecules that have an affinity to the gold surfaces of the nanoantennas used in detector prototypes. Self-assembled monolayers of thiol-containing molecules have been shown to have high stability and reproducibility on gold surfaces [15], which are often used in plasmonic devices. It is therefore prudent to focus on thiol-containing molecules that are commercially available or easily synthesizable. These requirements pose a challenge for high-throughput screening methods as the number of thiol compounds within large commercial databases is relatively low, with only around out of 18 million compounds in the eMolecules database and out of 8 million compounds from the MolPort database identified in Koczor-Benda et al. [10] being monothiols, respectively.
@@
An alternative solution for accelerating the discovery of promising molecules is generative deep learning, which in the past has been used for the property-driven design of functional organic molecules [16, 17, 18, 19, 20, 21]. Most proposed generative deep learning models use text-based or two-dimensional (2D) molecular representations [22, 23]. G-SchNet is a generative autoregressive deep neural network that has the advantage of being able to generate molecules in three-dimensional (3D) space [24]. Previous studies have shown that G-SchNet can be iteratively biased to generate molecules satisfying certain target properties. Westermayr et al. [16] used G-SchNet coupled with a neural network that predicts molecular quasiparticle energies [25] to bias G-SchNet prediction towards small fundamental gaps, low ionization potentials, or high electron affinities while conserving low synthetic complexity of the molecules. Gebauer et al. [17] developed conditional G-SchNet, which, in addition to structures, trains on electronic property and structural motif labels with which the generation can be conditioned.
@@
In this paper, we perform property-driven generative design of functional organic molecules for terahertz radiation detection using G-SchNet, driving the generative model to create novel molecules with high-frequency upconversion efficiency, affinity to gold surfaces, and synthetic accessibility. To increase the pool of candidates for this application, we train G-SchNet models on a dataset of around thiol-containing molecules and generate hundreds of thousands of monothiolated molecules by iterative biasing. We analyze chemical trends in the generated databases and identify functional groups that correlate with high upconversion intensity. Previously used ML predictors of the frequency upconversion efficiency based on molecular fingerprints [10] become unreliable as the property-driven generative biasing workflow explores novel molecules beyond the training dataset. We replace them with more transferable equivariant graph neural network (GNN) models that make use of the 3D molecular conformations that G-SchNet generates. Finally, highly spectroscopically active compounds are identified by generative design and further validated with quantum chemistry calculations and retrosynthetic route planning to identify promising, novel compounds for terahertz radiation detection.
@@ II.1 Generative machine learning
A training dataset of monothiolated molecules was compiled from the eMolecules [26] commercial molecular database, that was previously used by Koczor-Benda et al. [10]. The eMolecules database was first filtered for monothiols based on the corresponding SMARTS pattern. Charged molecules and duplicates were removed, resulting in molecules containing the following elements: hydrogen (H), boron (B), carbon (C), nitrogen (N), oxygen (O), fluorine (F), silicon (Si), phosphorus (P), sulfur (S), chlorine (Cl), selenium (Se), bromine (Br), tin (Sn), and iodine (I). In contrast to Koczor-Benda et al. [10], molecular size and number of rotatable bonds were not restricted, resulting in a larger pool of molecules. Initial 3D structures were created from Simplified Molecular Input Line Entry System (SMILES) strings [27] and relaxed with the MMFF94 Merck molecular force field [28] using the RDKit package [29]. To maximize chemical diversity, a Smooth Overlap of Atomic Positions (SOAP) [30] descriptor with a local region cut-off of , 4 radial basis functions, and a maximal degree of spherical harmonics of 3 was calculated for each molecule (resulting in features), using the DScribe package [31]. After singular value decomposition with 500 components, clusters were identified with -means clustering using the scikit-learn [32] library. For each cluster, the molecule closest to the cluster centre was selected. Molecules that had already been calculated in the THz database were removed (604 duplicates). Structure optimization was performed with the xTB software package using the GFN2-xTB parametrization [33], based on which the final database for the generative model was constructed.
@@
The schnetpack-gschnet [34, 35] package was used to train G-SchNet models on the aforementioned training database. Each G-SchNet model was trained using a SchNet [36] neural network with 128 features, 9 interaction blocks, a cut-off of and 25 centres for the radial basis expansion of distances. A learning rate of 0.0001 was used and 5 random atom placements per molecule per batch were drawn. For all trained G-SchNet models, data was split 80%/10%/10% for training, validation and testing, respectively. Approximately molecules were generated with each trained model, with a maximum molecular size of 60 atoms. Non-unique, disconnected, or invalid (incorrect valency) generated molecules were discarded. Molecules were filtered to only contain one thiol group, which can act as the linker to the gold nanoantenna in a THz radiation detector device. The number of molecules generated and remaining after filtering are summarized in the ESI (Table SI).
@@
The generation of molecules with desired properties was achieved by an iterative workflow similar to the one proposed by Westermayr et al. [16] Herein, in each iteration, the G-SchNet model is trained, molecules are generated, molecules are filtered with a property prediction model, and a new training dataset is built that contains the original and a subset of the novel generated molecules with selected properties above or below a certain threshold value. As a result, the molecule generation is iteratively biased towards molecules with desired properties. In each iteration, G-SchNet was trained (from scratch) with the modified dataset. The size of the training databases for each of the six biasing iterations is detailed in Table SII in the ESI.
In each iteration, molecules were selected according to two properties: the THz upconversion efficiency, predicted with a previously trained Kernel Ridge Regression (KRR) model [10], and the SCScore metric of synthetic complexity [37]. The upconversion efficiency figure of merit, P, is defined as the logarithm of the orientation-averaged upconversion intensity () summed over all vibrational frequencies in the 1–30 THz frequency window (1–1000 cm-1): [10]
Higher P values correspond to greater total frequency upconversion intensity of vibrations in the selected frequency range. A full definition of P can be found in the ESI (section S2). The SCScore neural network by Coley et al. [37] was trained on 12 million reactions from the Reaxys [38] database. The SCScore correlates with the number of reaction steps required to synthesize the molecule from reasonable starting materials and ranges between 1 and 5, where higher numbers indicate reduced synthesizability [37]. Canonical SMILES [27] representations of molecules generated using Open Babel [39] were used as input for the KRR predictor and the SCScore calculator. To simultaneously bias molecular generation towards large P (high THz upconversion efficiency) and low SCScore (S, low synthetic complexity) values, molecules with properties satisfying both and were appended to the training dataset for the subsequent training iteration, where and are the mean average and standard deviation, respectively, of property X.
@@
As reference data for the ML models, a database of about gold-thiolate molecules, available from Molecular Vibration Explorer [11], was used, henceforth referred to as the ‘THz database’. This database was originally compiled in Koczor-Benda et al. [10] and contains P values calculated with Kohn-Sham density functional theory (DFT) [40, 41], using the B3LYP [42, 43] hybrid generalized gradient approximation, the DFT-D3 [44] dispersion correction, the Karlsruhe basis set with split valence polarization (def2-SVP) [45], and a tight energy convergence threshold. To assess the accuracy of ML property predictors along the biasing iterations, additional reference calculations at the same level of theory were performed whereby the thiol group in each molecule was modified to a gold-thiolate group. The Gaussian16 [46] software package was used to run DFT calculations and analysis tools from Molecular Vibration Explorer [11] were used to calculate P values. The pretrained KRR model from Koczor-Benda et al. [10] was used to calculate P values; additionally, PaiNN [47] and MACE [48] equivariant GNN models were trained on the P values of the THz database. Full details of training and hyperparameter optimization, as well as learning curves, are provided in the ESI (Table SIII and SIV, Figures S1-S3).
@@ III.1 Analysis of generated molecules
The G-SchNet generative model is initially trained on the original dataset and used to generate novel and ”unbiased” molecules. A subset of the generated molecules is selected according to their predicted THz upconversion efficiency (high P value) and synthetic complexity (low SCScore) and added to the dataset. This process is repeated in six successive iterations during which properties of the generated molecules are driven towards the desired ranges (Figure 1a and b). Iterative biased generation of molecules successfully leads to molecules with higher P and lower SCScore in later iterations when compared to the training dataset (‘Train‘) and the unbiased initial generation (‘Unbiased‘). Further shifts in property values after iteration 5 were not significant and biasing was stopped after Iteration 6.
The composition of generated compounds differs significantly from the training set, as shown by the elemental composition of molecules in Fig. 1d. The differences are largest between the training set and the unbiased generated molecules, which highlights the fact that G-SchNet, without biasing or conditioning, does not fully reproduce the chemical features of its training set. This shortcoming has been previously observed by Westermayr et al. [16] and Gebauer [59]. This effect is more significant for models trained on diverse datasets featuring many elements and molecular sizes than for models trained on small and simple molecules (such as QM9 [60, 61]). The unbiased generated molecules feature a significantly reduced proportion of hydrogen atoms compared to the training dataset, which suggests increased numbers of unsaturated bonds and heteroatomic groups. The proportion of hydrogen atoms slightly increases through the subsequent biased iterations. Nitrogen atoms also become more prevalent in generated sets, while the proportion of carbon and fluorine atoms decreases. There is a shift of the size distribution of molecules to smaller values, as shown in Fig. S4a. While unbiased generation creates significant numbers of molecules with 30–60 atoms, generated molecules in later iterations have, on average, about 20 atoms. A significant number of molecules generated by the unbiased model have an SCScore above 4 (Fig. 1b, which was also observed by Westermayr et al. [16]. We note that all training molecules are commercially available so the SCScore metric does not fully reflect their accessibility but rather was used as an indicative metric by which we filter out generated molecules that are overly complex. For the most promising generated candidate molecules, we perform comprehensive retrosynthetic planning analysis to assess their synthesizability more accurately (vide infra).

### v2

@@ Abstract
Plasmonic nanocavities are molecule-nanoparticle junctions that offer a promising approach to upconvert terahertz radiation into visible or near-infrared light, enabling nanoscale detection at room temperature. However, the identification of molecules with strong terahertz-to-visible frequency upconversion efficiency is limited by the availability of suitable compounds in commercial databases. Here, we employ the generative autoregressive deep neural network, G-SchNet, to perform property-driven design of novel monothiolated molecules tailored for terahertz radiation detection. To design functional organic molecules, we iteratively bias G-SchNet to drive molecular generation towards highly active and synthesizable molecules based on machine learning-based property predictors, including molecular fingerprints and state-of-the-art neural networks. We study the reliability of these property predictors for generated molecules and analyze the chemical space and properties of generated molecules to identify trends in activity. Finally, we filter generated molecules and plan retrosynthetic routes from commercially available reactants to identify promising novel compounds and their most active vibrational modes in terahertz-to-visible upconversion.
@@ I Introduction
Terahertz (THz) radiation has applications in numerous fields, including medical diagnostics, security screening, communications, and astronomy [1, 2]. Historically, the development of both powerful and affordable light sources, and efficient THz detectors, has been technologically challenging.
@@
Nanoscale, room-temperature detection of terahertz and mid-infrared radiation is enabled by molecular optomechanical devices utilizing the enhancement of electronic fields in plasmonic nanocavities to convert terahertz radiation into visible or near-infrared light [3, 4]. These nanocavities can be assembled on silicon-based photonic integrated circuits [5], opening possibilities for low-cost fabrication and multiplexed detection. To enhance the light-matter interaction, molecules are typically placed between two metallic nanoantennas [4, 6, 7]. One of the two antennas focuses terahertz radiation at the design frequency over the molecular sample volume to enhance the absorption of terahertz radiation via the surface-enhanced infrared absorption [8] mechanism. The second optical antenna confines visible or near-infrared light to volumes below 100 , which induces surface-enhanced Raman scattering [9] of molecules within the plasmonic nanocavity. Absorption of THz radiation by molecules within the nanocavity results in the vibrational excitation of a specific normal mode, which leads to an increase in the measured Raman anti-Stokes intensity of the same normal mode, similar to resonant sum-frequency generation spectroscopy [10]. For centrosymmetric molecules, simultaneous activity in absorption and Raman scattering is not possible. Even in asymmetric molecules, it is rare to have vibrational modes that can efficiently upconvert the THz radiation signal, as this requires a large change in both electronic dipole moment and in polarizability along the vibrational mode. Vibrational modes of organic molecules in the THz frequency range are often delocalized across several functional groups or across molecules, which makes it challenging to use chemical intuition to suggest promising candidates or define molecular design rules. This makes it necessary to use quantum chemical calculations in connection with computational screening or design approaches to identify good candidate molecules and their active vibrational modes [11, 12]. Such computational predictions motivate more detailed experimental investigations for the fabrication and application of new molecular optomechanical devices [5].
@@
Machine learning (ML) methods can facilitate the design and discovery of new functional materials by enabling the fast computational screening of large structural databases [13, 14, 15]. ML-based screening has previously been used to identify promising candidates for THz radiation detection from commercially available compound databases [11]. However, a drawback of this approach was that there was a limited search pool of molecules that have an affinity to the gold surfaces of the nanoantennas used in detector prototypes. Self-assembled monolayers of thiol-containing molecules have been shown to have high stability and reproducibility on gold surfaces [16], which are often used in plasmonic devices. It is therefore prudent to focus on thiol-containing molecules that are commercially available or easily synthesizable. These requirements pose a challenge for high-throughput screening methods as the number of thiol compounds within large commercial databases is relatively low, with only around out of more than 20 million compounds in the eMolecules database and out of 8 million compounds from the MolPort database identified in Koczor-Benda et al. [11] being monothiols, respectively.
@@
An alternative solution for accelerating the discovery of promising molecules is generative deep learning, which in the past has been used for the property-driven design of functional organic molecules [17, 18, 19, 20, 21, 22]. Most proposed generative deep learning models use text-based or two-dimensional (2D) molecular representations [23, 24]. G-SchNet is a generative autoregressive deep neural network that has the advantage of being able to generate molecules in three-dimensional (3D) space [25]. Previous studies have shown that G-SchNet can be iteratively biased to generate molecules satisfying certain target properties. Westermayr et al. [17] coupled G-SchNet with a neural network that predicts molecular quasiparticle energies [26] to bias molecular generation towards small fundamental gaps, low ionization potentials, or high electron affinities, while conserving low synthetic complexity of the molecules. Gebauer et al. [18] developed conditional G-SchNet, which, in addition to structures, trains on electronic property and structural motif labels in order to condition molecular generation.
@@
In this paper, we perform property-driven generative design of functional organic molecules for THz radiation detection using G-SchNet by driving the generative model to create novel molecules with high frequency-upconversion efficiency, affinity to gold surfaces, and synthetic accessibility. To predict the upconversion properties of molecules, we use the target property P introduced in Koczor-Benda et al. [11], which is based on the total spectral intensity in a wide frequency window (30–1000 ) relevant for THz and mid-infrared applications. To increase the pool of candidates for this application, we train G-SchNet models on a dataset of around thiol-containing molecules and generate hundreds of thousands of monothiolated molecules by iterative biasing. We analyze chemical trends in the generated databases and identify functional groups that correlate with high upconversion intensity. Previously used ML predictors of the frequency upconversion efficiency based on molecular fingerprints [11] become unreliable as the property-driven generative biasing workflow explores novel molecules beyond the training dataset. We replace them with more transferable equivariant graph neural network (GNN) models that make use of the 3D molecular conformations that G-SchNet generates. To train these models, we use calculations based on density functional theory (DFT) for P values contained in Molecular Vibration Explorer [12], which are available for around 2800 gold-thiolate molecules, and extend this database with new DFT calculations on generated molecules. Finally, highly spectroscopically active compounds are identified by generative design and further validated with quantum chemistry calculations and retrosynthetic route planning to identify promising, novel compounds for THz radiation detection.
@@ II.1 Generative machine learning
A training dataset of monothiolated molecules was compiled from the eMolecules [27] commercial molecular database, that was previously used by Koczor-Benda et al. [11]. This database contains over 20 million readily available or custom synthesized compounds from over 15 suppliers, aimed mainly at drug discovery applications [27]. This training dataset was selected to ensure that the generative model creates molecules that are chemically similar to known synthesizable compounds, thus facilitating the search for viable candidates. The eMolecules database was first filtered for monothiols based on the corresponding SMARTS pattern. Charged molecules and duplicates were removed, resulting in an initial pool of molecules containing the following elements: hydrogen (H), boron (B), carbon (C), nitrogen (N), oxygen (O), fluorine (F), silicon (Si), phosphorus (P), sulfur (S), chlorine (Cl), selenium (Se), bromine (Br), tin (Sn), and iodine (I). In contrast to Koczor-Benda et al. [11], molecular size and number of rotatable bonds were not restricted, resulting in a larger pool of molecules. Initial 3D structures for the unique monothiolated molecules were created from Simplified Molecular Input Line Entry System (SMILES) strings [28] and relaxed with the MMFF94 Merck molecular force field [29] using the RDKit package [30]. To maximize chemical diversity, a Smooth Overlap of Atomic Positions (SOAP) [31] descriptor with a local region cut-off of , 4 radial basis functions, and a maximal degree of spherical harmonics of 3 was calculated for each molecule (resulting in features), using the DScribe package [32]. After singular value decomposition with 500 components, clusters were identified with -means clustering using the scikit-learn [33] library. For each cluster, the molecule closest to the cluster center was selected. Molecules that had already been calculated in the THz database were removed (604 duplicates), resulting in the final training set of molecules. Structure optimization was performed with the xTB software package using the GFN2-xTB parametrization [34], based on which the final database of 3D geometries for the generative model was constructed. For a discussion on using the computationally less expensive xTB method instead of commonly used DFT for structure optimisation, and a comparison of generated unrelaxed and relaxed structures see section S10 in the ESI.
@@
The schnetpack-gschnet [35, 36] package was used to train G-SchNet models on the aforementioned training database. Each G-SchNet model was trained using a SchNet [37] neural network with 128 features, 9 interaction blocks, a cut-off of and 25 centers for the radial basis expansion of distances. A learning rate of 0.0001 was used and 5 random atom placements per molecule per batch were drawn. For all trained G-SchNet models, data were randomly split (as implemented within schnetpack-gschnet) 80%/10%/10% for training, validation and testing, respectively. Approximately molecules were generated with each trained model, with a maximum molecular size of 60 atoms. Non-unique, disconnected, or invalid (incorrect valency) generated molecules were discarded. Molecules were filtered to only contain one thiol group, which can act as the linker to the gold nanoantenna in a THz radiation detector device. The number of molecules generated and remaining after filtering are summarized in the ESI (Table SI).
@@
The generation of molecules with desired properties was achieved by an iterative workflow similar to the one proposed by Westermayr et al. [17] Herein, in each iteration, the G-SchNet model is trained, molecules are generated, molecules are filtered with a property prediction model, and a new training dataset is built that contains the original and a subset of the novel generated molecules with selected properties above or below a certain threshold value. As a result, molecular generation is iteratively biased towards molecules with desired properties. In each iteration, G-SchNet was trained (from scratch) with the modified dataset. The sizes of the training databases for each of the six biasing iterations are detailed in Table SII in the ESI.
In each iteration, molecules were selected according to two properties: THz upconversion efficiency, predicted with a previously trained Kernel Ridge Regression (KRR) model [11], and the SCScore metric of synthetic complexity [38]. The upconversion efficiency figure of merit, P, is defined as the logarithm of the orientation-averaged upconversion intensity () summed over all vibrational frequencies in the 1–30 frequency window (30–1000 ): [11]
Higher P values correspond to greater total frequency upconversion intensity of vibrations in the selected frequency range. is based on the absorption and Raman scattering intensities of vibrational mode (a full definition of can be found in section S2 of the ESI). was calculated using DFT for a simplified model of the molecule-metal interface in Koczor-Benda et al. [11] resulting in the Gold database of Molecular Vibration Explorer [12], and used as training data for the KRR model [11]. We also discuss the details of these DFT calculations in the next section.
The SCScore neural network by Coley et al. [38] was trained on 12 million reactions from the Reaxys [39] database. The SCScore correlates with the number of reaction steps required to synthesize the molecule from reasonable starting materials and ranges between 1 and 5, where higher numbers indicate reduced synthesizability [38]. Canonical SMILES [28] representations of molecules generated using Open Babel [40] were used as input for the KRR predictor and the SCScore calculator. To simultaneously bias molecular generation towards large P (high THz upconversion efficiency) and low SCScore (S, low synthetic complexity) values, molecules with properties satisfying both and were appended to the training dataset for the subsequent training iteration, where and are the mean average and standard deviation, respectively, of property X.
@@
As reference data for the ML models, a database of about gold-thiolate molecules, available from Molecular Vibration Explorer [12], was used, henceforth referred to as the ‘THz database’. This database was originally compiled in Koczor-Benda et al. [11] and contains P values calculated with Kohn-Sham DFT [41, 42], using the B3LYP [43, 44] hybrid generalized gradient approximation, the DFT-D3 [45] dispersion correction, the Karlsruhe basis set with split valence polarization (def2-SVP) [46], and a tight energy convergence threshold. The molecules were modeled as gold-thiolates to consider the most immediate chemical effects of the metal-molecule interface. This choice of modeling was also validated against surface-enhanced Raman spectroscopy measurements in Griffiths et al. [47], Boehmke Amoruso et al. [48], and Wright et al. [49]. Koczor-Benda et al. [11] validated the computational approach in detail against Raman and infrared measurements for powder, solution and nanoparticle-on-mirror constructs of a set of test molecules and found that individual spectral features as well as surface-enhanced Raman spectroscopy intensities integrated over a wide spectral window correlate well with measurements. However, for an accurate modeling of low-frequency vibrational features (below 200 ), considering the metal facets as well as molecule-molecule interactions becomes necessary [48], which increases computational costs. To enable a fast computational assessment of a large number of molecules, and benefit from the existing, openly available database we follow the approach of Koczor-Benda et al. [11]. To assess the accuracy of ML property predictors along the biasing iterations, additional reference calculations at the same level of theory were performed whereby the thiol group in each molecule was modified to a gold-thiolate group. The Gaussian16 [50] software package was used to run DFT calculations and analysis tools from Molecular Vibration Explorer [12] were used to calculate P values. The pretrained KRR model from Koczor-Benda et al. [11] was used to predict P values; additionally, PaiNN [51] and MACE [52] equivariant GNN models were trained on the P values of the DFT-optimized structures of the THz database. Full details of training and hyperparameter optimization, as well as learning curves, are provided in the ESI (Table SIII and SIV, Figures S1-S3). The PaiNN and MACE predictions of P values are based on the unrelaxed 3D structures of generated molecules, following Westermayr et al. [17]. Section 10 of the ESI discusses the effect of using unrelaxed structures on the predicted P values for a subset of generated molecules.
@@ III.1 Analysis of generated molecules
The G-SchNet generative model is initially trained on the original dataset and used to generate novel and ‘unbiased’ molecules. A subset of the generated molecules is selected according to their predicted THz upconversion efficiency (high P value) and synthetic complexity (low SCScore) and added to the dataset. This process is repeated in six successive iterations during which properties of the generated molecules are driven towards the desired ranges (Figure 1a and b). Iterative biased generation of molecules successfully leads to molecules with higher P and lower SCScore in later iterations when compared to the training dataset (‘Train’) and the unbiased initial generation (‘Unbiased’). Further shifts in property values after iteration 5 were not significant and biasing was stopped after Iteration 6.
The composition of generated compounds differs significantly from the training set, as shown by the elemental composition of molecules in Fig. 1d. The differences are largest between the training set and the unbiased generated molecules, which highlights the fact that G-SchNet, without biasing or conditioning, does not fully reproduce the chemical features of its training set. This shortcoming has been previously observed by Westermayr et al. [17] and Gebauer [63]. This effect is more significant for models trained on diverse datasets featuring many elements and molecular sizes than for models trained on small and simple molecules (such as QM9 [64, 65]). The unbiased generated molecules feature a significantly reduced proportion of hydrogen atoms compared to the training dataset, which suggests increased numbers of unsaturated bonds and heteroatomic groups. The proportion of hydrogen atoms slightly increases through the subsequent biased iterations. Nitrogen atoms also become more prevalent in generated sets, while the proportion of carbon and fluorine atoms decreases. There is a shift of the size distribution of molecules to smaller values, as shown in Fig. S4a. While unbiased generation creates significant numbers of molecules with 30–60 atoms, generated molecules in later iterations have, on average, about 20 atoms. A significant number of molecules generated by the unbiased model have an SCScore above 4 (Fig. 1b), which was also observed by Westermayr et al. [17]. We note that all training molecules are commercially available so the SCScore metric does not fully reflect their accessibility but rather was used as an indicative metric by which we filter out generated molecules that are overly complex. For the most promising generated candidate molecules, we perform comprehensive retrosynthetic planning analysis to assess their synthesizability more accurately (vide infra).